# Efficient Bayesian Updates for Deep Learning via Laplace Approximations

### v2

@@ Abstract
Since training deep neural networks takes significant computational resources, extending the training dataset with new data is difficult, as it typically requires complete retraining. Moreover, specific applications do not allow costly retraining due to time or computational constraints. We address this issue by proposing a novel Bayesian update method for deep neural networks by using a last-layer Laplace approximation. Concretely, we leverage second-order optimization techniques on the Gaussian posterior distribution of a Laplace approximation, computing the inverse Hessian matrix in closed form. This way, our method allows for fast and effective updates upon the arrival of new data in a stationary setting. A large-scale evaluation study across different data modalities confirms that our updates are a fast and competitive alternative to costly retraining. Furthermore, we demonstrate its applicability in a deep active learning scenario by using our update to improve existing selection strategies.
@@ 1 Introduction
Training deep neural networks (DNNs) often demands significant time and computational resources. Moreover, when extending a dataset with new data, DNNs require complete retraining, which involves both the new and previously used data to prevent issues such as catastrophic forgetting [42]. Although improved generalization performance typically justifies exhaustive retraining, it may not always be feasible immediately after new data arrives. Key scenarios where retraining is beneficial but impractical include: (1) When data arrives sequentially and an immediate update is necessary, e.g., in active learning (AL) [48] or when working on data streams [46], (2) when only limited computational resources are available, e.g., training on embedded hardware [51], and (3) when there are privacy concerns, which restrict sharing new data with distributed computing units [51]. In these scenarios, it is crucial to devise methods that allow DNNs to incorporate new data fast and effectively.
While numerous studies developed methods to update DNNs effectively [19, 42, 50], these methods typically assume a non-stationary setting in which the respective data distribution changes over time [52]. As a result, these methods often focus on situations where new data arrives in relatively large amounts [42], neglecting scenarios where an immediate update with a small number of instances might be beneficial. In contrast, we assume a stationary setting with the goal to immediately update the model upon the arrival of new data, even in the case of a single data point.
As an example scenario, we consider AL throughout this article. A popular approach in AL is to use a one-step-look-ahead to select instances that significantly change model predictions [44, 50]. This approach assesses how model predictions would change, when including a specific instance in the labeled pool and updating the model. Consequently, it requires a highly efficient update method, as it involves updates for all candidates of a large dataset. To make this feasible, several AL selection strategies utilize a Bayesian neural network (BNN) and rely on Monte Carlo (MC)-based updates [50]. However, while this update is fast, it falls short in performance compared to a retrained model (cf. Fig. 1). In such a scenario, a fast and effective update would enhance these strategies.
In this article, we propose a novel Bayesian update method. Like [50, 42, 21], we consider BNNs since they are particularly well-suited for our setting. The posterior distribution can be updated in a theoretically sound way by treating it as the prior distribution when new data becomes available. However, instead of relying on MC-based BNNs (e.g., MC-Dropout), we transform an arbitrary DNN into a BNN by employing a last-layer Laplace approximation (LA) [7], giving us a closed-form expression for the posterior distribution. Then, we leverage second-order optimization techniques on the Gaussian posterior distribution of the LA. To ensure low computational complexity, we compute the required inverse Hessian in closed form. The resulting update is both fast – it can be used in the previously described scenarios – and effective – its performance closely aligns with retraining. Moreover, our update can be combined with more recently developed BNNs such as Spectral Normalized Neural Gaussian Process (SNGP) [30].
Extensive studies across different data modalities, including image and text datasets, demonstrate that our updates outperform the typically employed MC-based ones [21, 50] in terms of speed and performance. Further, to demonstrate the applicability in one of the previously mentioned scenarios, we revisit AL and propose a simple and effective idea to improve existing selection strategies by immediately making use of acquired labels.
Figure 1: Comparison of different BNNs on the two moons dataset. (a) The original model resulting from training on the original dataset. (b) The typical MC-based update applied to the original model [21, 50]. (c) Our update applied to the original model. (d) The model resulting from retraining on both the original and the new dataset.
Contributions We propose a novel update method for DNNs by employing a last-layer LA and second-order optimization techniques, suitable for scenarios where data arrives in small quantities, and immediate model updates are required. We conduct a comprehensive evaluation of our update method across different data modalities, demonstrating superior performance and speed compared to the MC-based updates. We propose a simple framework to improve existing AL strategies employing our updates.
@@ 2 Related Work
Similar to our setting, continual learning [8] updates models by exclusively training with data from a new task, addressing the challenge of retaining knowledge from previously learned tasks. Popular techniques [42, 19] use conventional first-order optimization methods, incorporating a regularization term to counteract catastrophic forgetting. More specifically, [42] and [19] derive a regularization term from an LA that penalizes large deviations from prior knowledge. Ebrahimi et al. [11] exploit uncertainty estimates of BNNs to dynamically adjust learning rates during training. Unlike our method, these approaches require training over multiple epochs, preventing immediate updates. Additionally, they require large amounts of new data (thousands of data points) per task, whereas our update method is designed for smaller datasets, ranging from single to hundreds of data points. More closely related to our work is online learning [17], which aims to sequentially and efficiently update models from incoming data streams. Traditional approaches often focus on linear [57, 6] or shallow [22, 45] models with maximum-margin classification. However, applying online learning to DNNs remains difficult due to issues such as convergence, vanishing gradients, and large model sizes [46, 27]. To address these challenges, Sahoo et al. [46] proposed a method that modifies a DNN’s architecture to facilitate updates. We argue that this approach is restrictive in state-of-the-art settings, given the increasing reliance on pretrained architectures. Most similar to our setting is the work on Bayesian online inference by Kirsch et al. [21]. The core idea is to sample hypotheses, e.g., via MC-Dropout, from the posterior distribution of a BNN and weight their importance according to the respective likelihoods for sequentially arriving data. We refer to these types of updates as MC-based updates. The empirical results raised concerns regarding the applicability of such MC updates in high-dimensional parameter spaces.
BNNs [53, 12] induce a prior distribution over their parameters, i.e., weights, and learn a posterior distribution given training data. Predictions are made by marginalizing this posterior. BNN types differ mainly in their probabilistic model and sampling from the posterior distribution [18]. MC-Dropout [13], one of the most prominent BNN types, is a regularization technique performed during training. Using dropout during evaluation, called MC-Dropout, we obtain a distribution for the predictions corresponding to a variational distribution in the parameter space. Due to its simplicity and training efficiency, MC-Dropout is often used for comparison. However, its inference is inefficient and predictions may not properly represent uncertainty estimates [37]. Deep ensembles [25] consist of multiple DNNs. Combined with regularization, these DNNs are samples (different modes) of the parameters’ posterior distribution. Ensembles typically provide better uncertainty estimates than MC-Dropout but require more computational capacity during training [37]. A BNN obtained via LA [43] is computationally more efficient than MC-Dropout and deep ensembles. It specifies an approximate Gaussian posterior distribution, where the maximum a posteriori (MAP) estimate defines the mean and the inverse of the negative log posterior’s Hessian at the MAP estimate corresponds to the covariance matrix. As computing this Hessian is expensive for large DNNs, LA is often used only in the last layer [7]. SNGP [30] combines LA with random Fourier features [38] and spectral normalization [34] to approximate a Gaussian process.
AL poses significant challenges in practice due to the computational cost associated with required retraining per cycle. The emergence of one-step-look-ahead strategies [44, 50, 56] further highlights the importance of efficient model updates. Notably, BEMPS [50] evaluates all possible candidates (i.e., instances) in a dataset to select those that are anticipated to alter the model’s predictions substantially. However, this strategy utilizes an MC-based BNN with deep ensembles, relying on MC-based updates. They assume that all hypotheses are equally likely to explain the data before an update step. While this update is fast, its performance is inferior to that of a retrained model.
@@ 3 Fast Bayesian Updates for Deep Neural Networks
In this section, we present our new Bayesian update method. First, we introduce the general concept of Bayesian updates and the commonly applied MC-based Bayesian updates [21, 50]. Afterward, we propose our novel method focusing on an efficient update of the Gaussian posterior distribution via last-layer LAs. For an introduction to LA, we refer to [7].
@@ 3.1 Bayesian Updates
We focus on classification problems with instance space $\mathcal{X}$ and label space $\mathcal{Y}=\{0,\dots,K-1\}$. The primary goal in our setting is to efficiently incorporate the information of new instance-label pairs ${\mathcal{D}^{\oplus}=\{({\boldsymbol{x}}_{n},y_{n})\}_{n=1}^{N}\subset\mathcal{X}\times\mathcal{Y}}$ into a BNN trained on dataset $\mathcal{D}\subset\mathcal{X}\times\mathcal{Y}$. Retraining the entire network on the extended dataset $\mathcal{D}\cup\mathcal{D}^{\oplus}$ results in high computational cost for a large dataset $\mathcal{D}$. Conversely, using the new data solely can cause catastrophic forgetting [42].
For this purpose, we employ BNNs [12] with Bayesian updates [35] as an efficient alternative to retraining. The main idea of BNNs is to estimate posterior distribution $p(\boldsymbol{\omega}|\mathcal{D})$ over the parameters $\boldsymbol{\omega}\in\Omega$ given the observed training data $\mathcal{D}$ using Bayes’ theorem. The obtained posterior distribution over the parameters can then be used to specify the predictive distribution over a new instance’s class membership via marginalization:
@@
Thereby, the likelihood $p(y|{\boldsymbol{x}},\boldsymbol{\omega})=[\mathrm{softmax}(f^{\boldsymbol{\omega}}({\boldsymbol{x}}))]_{y}$ denotes the probabilistic output of a DNN with parameters $\boldsymbol{\omega}$, where $f^{\boldsymbol{\omega}}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ is a function outputting class-wise logits. [^1] We denote the $i$-th element of a vector ${\boldsymbol{b}}$ as $[{\boldsymbol{b}}]_{i}=b_{i}$.
The formulation in Eq. (1) provides a theoretically sound way to obtain updated predictions. In particular, this is because the probabilistic outputs $p(y|{\boldsymbol{x}},\boldsymbol{\omega})$ do not directly depend on the training data $\mathcal{D}$. Consequently, to obtain an updated predictive distribution, we do not need to update the parameters $\boldsymbol{\omega}$ directly but only the posterior distribution $p(\boldsymbol{\omega}|\mathcal{D})$. The updated posterior distribution $p(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ is found through Bayes’ theorem, where the current posterior distribution $p(\boldsymbol{\omega}|\mathcal{D})$ is considered the prior and multiplied with the likelihood $p(y|{\boldsymbol{x}},\boldsymbol{\omega})$ per instance-label pair $({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}$. As instances in $\mathcal{D}$ and $\mathcal{D}^{\oplus}$ are assumed to be independently distributed, we can simplify the likelihood and reformulate the parameter distribution as follows [^2] We denote $p(y_{1},\dots,y_{N}\mid{\boldsymbol{x}}_{1},\dots,{\boldsymbol{x}}_{N},\boldsymbol{\omega})$ with $\mathcal{D}=\{({\boldsymbol{x}}_{n},y_{n})\}_{n=1}^{N}$ as $p(\mathcal{D}|\boldsymbol{\omega})$.:
$p(\boldsymbol{\omega}|\mathcal{D}^{\oplus},\mathcal{D})\propto p(\boldsymbol{\omega}|\mathcal{D})p(\mathcal{D}^{\oplus}|\mathcal{D},\boldsymbol{\omega})\overset{\mathrm{i.i.d.}}{=}p(\boldsymbol{\omega}|\mathcal{D})p(\mathcal{D}^{\oplus}|\boldsymbol{\omega})=p(\boldsymbol{\omega}|\mathcal{D})\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}p(y|{\boldsymbol{x}},\boldsymbol{\omega}).$ (2)
We refer to Eq. (2) as the Bayesian update.
@@
The most common realization [21, 50] of this update is through MC-based BNNs, such as MC-Dropout and deep ensembles. These BNNs rely on samples (or hypotheses) $\boldsymbol{\omega}_{1},\dots,\boldsymbol{\omega}_{M}$ drawn from an approximate posterior $q(\boldsymbol{\omega}|\mathcal{D})$ Research [56, 50] assumes that all hypotheses are equally likely to explain the observed data and have the same probability before updating. By updating the posterior distribution through Eq. (2), they weigh more likely hypotheses given the new data higher. We refer to these as MC-based updates with a formal definition given in Appendix A. Figure 2 illustrates this concept where different hypotheses $\boldsymbol{\omega}_{1},\dots,\boldsymbol{\omega}_{M}\sim q(\boldsymbol{\omega}|\mathcal{D})$ are shown. Each hypothesis represents a possible true solution for the learning task (white instances). When new data (green instances) arrives, we weigh each hypothesis by its likelihood of explaining the new data and obtain an updated prediction without retraining. This results in an updated predictive distribution, as seen in bold in Fig. 2 (right).

### v3

# Efficient Bayesian Updates for Deep Active Learning via Laplace Approximations

@@ Abstract
Deep active learning (AL) selects batches of instances for annotation to avoid retraining deep neural networks (DNNs) after each new label. Employing a naive top-$b$ selection can result in a batch of redundant (similar) instances. To address this, various AL strategies employ clustering techniques that ensure diversity within a batch. We approach this issue by substituting the costly retraining with an efficient Bayesian update. Our proposed update represents a second-order optimization step using the Gaussian posterior from a last-layer Laplace approximation. Thereby, we achieve low computational complexity by computing the inverse Hessian in closed form. We demonstrate that in typical AL settings, our update closely approximates retraining while being considerably faster. Leveraging our update, we introduce a new framework for batch selection through sequential construction, updating the DNN after each label acquisition. Furthermore, we incorporate our update into a look-ahead selection strategy as a feasible upper baseline approximating optimal batch selection. Our results highlight the potential of efficient updates to advance deep AL research.
@@ 1 Introduction
Active Learning (AL) sequentially selects instances for annotation by human experts, aiming to maximize model performance while minimizing labeling efforts. When combined with deep neural networks (DNNs), AL typically selects instances in batches rather than one at a time. The reason for this is that retraining DNNs after each label acquisition is computationally expensive, and delay can lead to additional costs since annotators’ time is valuable [21].
In a naive top-$b$ batch selection, a batch of $b$ instances with the highest scores is chosen based on an informativeness measure. However, when many similar instances are present, this approach can result in significant redundancy within the batch (similar instances have a similar score). Many selection strategies have been developed to replace this naive selection [16]. These strategies often employ clustering techniques to ensure diverse batches, ensuring that instances within a batch are dissimilar to one another [16]. While effective in reducing redundancy, clustering does not guarantee optimal selection due to its heuristic motivation.
Orthogonal to these strategies, we explore the concept of efficient “retraining” in deep AL. If retraining were computationally feasible, researchers could place greater emphasis on the development of theoretically sound informativeness measures instead of using heuristic clustering approaches to ensure diversity. Additionally, selection strategies that aim to maximize future performance–strategies that have been shown to be near-optimal in traditional AL [35]–could be made feasible with DNNs. Therefore, we examine the concept of updating DNNs through a single optimization step as a proxy for retraining and explore its potential to enhance the AL process.
To underscore the requirements of such an update, consider strategies designed to maximize future performance. Typically, these use a look-ahead to select instances that significantly change model predictions. Specifically, they examine how adding unlabeled instances to the labeled pool and retraining the model affects predictions [35]. However, with a large number of unlabeled instances and the costly retraining process of DNNs, this approach becomes infeasible in deep AL. Therefore, instead of retraining, a highly efficient update method is required. The only work to realize this with DNNs is by [40], which employ an ensemble of DNNs combined with Monte Carlo (MC) updates via Bayes’ theorem. Although this update makes a look-ahead feasible, it remains suboptimal for several reasons: (i) the update requires an ensemble of DNNs, making the actual retraining time and memory demands inefficient; (ii) the update does not accurately reflect the performance of full retraining; and (iii) the updating process becomes inefficient with an increasing number of ensemble members.
In this article, we propose an efficient update method for DNNs in the context of AL for classification. Specifically, we transform an arbitrary DNN into a Bayesian neural network (BNN) by employing a last-layer Laplace approximation (LA) [8]. While the closed-form expression of the posterior allows us to leverage second-order optimization techniques, we ensure low computational complexity by computing the required inverse Hessian analytically. Unlike the MC-based update used in [40], our approach does not require an ensemble of DNNs, making it easily applicable and both memory- and training-efficient [8]. Additionally, as we utilize a single DNN, we can leverage pretrained foundation models [31], which are an essential part of modern AL strategies [16, 15]. The resulting update is fast and closely matches the performance of full retraining. Extensive studies across different data modalities demonstrate that our updates outperform the typically employed MC-based ones [40] in terms of speed and performance. Furthermore, we examine the proposed update in two distinct AL scenarios:
  1. 1.
    Enhancing Existing Strategies with Immediate Label Utilization: We propose a simple framework to improve existing strategies by immediately making use of acquired labels through the proposed updates. Rather than selecting the top-$b$ highest-scoring instances simultaneously, we iteratively select the highest-scoring instance $b$ times but update the model between each selection. This simple strategy, which approximates single-instance AL during batch construction, performs surprisingly well, outperforming naive top-$b$ selection as well as selection strategies that employ clustering.
  2. 2.
    Optimal AL with Look-Ahead Selection: We investigate the potential of our update with a look-ahead selection strategy in an optimal AL setting. Specifically, we approximate an optimal selection strategy that maximizes future performance. Instead of retraining, we ensure computational feasibility by employing our update. The resulting strategy outperforms all competitors, showcasing that currently employed selection strategies have much potential for improvement.
Summary of Contributions Efficient DNN Update: We propose an efficient update method for DNNs that employs a Laplace approximation and second-order optimization techniques. We enable low computational complexity through closed-form computation of the inverse Hessian. Comprehensive Evaluation: We perform an extensive evaluation across data modalities, demonstrating that our update outperforms MC-based updates in both speed and accuracy. Immediate Label Utilization: We develop a simple framework that employs our update to immediately incorporate acquired labels, improving existing selection strategies by updating the model during batch construction. Optimal AL with Look-Ahead: We study our update in an optimal AL setting, making a near-optimal selection strategy as an upper baseline computationally feasible.
@@ 2 Related Work
Pool-based deep AL strategies can be divided into three types. Uncertainty-based methods, such as margin sampling [3, 38] and BALD [13], assume that instances with high predictive uncertainty are most informative. When selecting batches, these strategies pick the top‑$b$ highest-scoring instances, leading to redundancy. In contrast, diversity-based approaches like Core-Set [37] aim to select a diverse batch of instances by considering the feature representations of labeled and unlabeled data. In practice, hybrid strategies, a combination of both types, have been shown to work well. BatchBALD [19] extends BALD by reducing redundant information within a batch. Badge [1] selects instances with high gradient norms and ensures diversity by employing $k$-means++ in the gradient space. Typiclust [16] replaces the notion of uncertainty with typicality and selects typical instances from clusters obtained through $k$-means.
Look-ahead strategies [35] remain underexplored in deep AL. They aim to select instances expected to improve the model’s performance the most by retraining for all possible candidate instances. In non-deep settings, such approaches have been shown to achieve near-optimal selection [35] while offering convergence guarantees [46]. However, adapting these strategies to deep AL is challenging due to the computational cost of retraining. To the best of our knowledge, BEMPS [40] is the only strategy to implement a look-ahead mechanism in deep AL. They employ deep ensembles and MC-based updates via Bayes’ theorem (see Sec.˜3). while this update is computationally efficient, we show that its performance falls short compared to full model retraining.
Continual learning [9] updates models by only using new data, retaining knowledge from the previous one. Related techniques [18, 34] use conventional first-order optimization methods, deriving regularization terms from an LA that penalizes large deviations from prior knowledge. Unlike our method, these approaches require training over multiple epochs for the regularization term to have an impact. Used as an update (i.e., single optimization step), these strategies simplify to a first-order gradient step. Additionally, they assume large amounts of new data per task (thousands of instances), whereas our update method is designed batch construction in AL, ranging from a single to hundreds of instances. For example, a typical benchmark is to extend a dataset with a task consisting of all instance-label pairs of a new class (ca. 5,000 in MNIST).
@@
More closely related to our work is online learning [17], which aims to sequentially and efficiently update models from incoming data streams. Traditional approaches often focus on linear [47] or shallow [22] models with maximum-margin classification. However, applying online learning to DNNs remains difficult due to issues such as convergence, vanishing gradients, and large model sizes [36, 44]. Sahoo et al. [36] proposed a method that modifies a DNN’s architecture to facilitate updates. We argue that this is restrictive in state-of-the-art settings, given the increasing reliance on pretrained foundation models [10, 31]. Recently, [20] proposed Bayesian online inference, which is also used in [40]. This method samples hypotheses (e.g., via MC-Dropout) from a BNN’s posterior and weights them by the likelihood for new arriving data. However, the empirical results raise concerns about its feasibility in high-dimensional parameter spaces. We refer to these as MC-based updates.
BNNs [41] induce a prior distribution on parameters of a DNN and learn a posterior distribution given data. MC-Dropout [12] uses dropout to obtain a distribution over predictions. While it is simple to use, its inference is inefficient, and it provides suboptimal uncertainty estimates [32]. Deep ensembles [24] are known for their superior uncertainty estimates but are train and memory inefficient [32]. LAs [8] approximate the posterior as a Gaussian, with the MAP estimate as the mean and the inverse Hessian as the covariance. As computing this Hessian is expensive for large DNNs, LA is often used only in the last layer [8].
Finally, we consider related approaches focusing on the efficiency of AL. Prior work [6] shows that smaller more efficient proxy models can be used for AL selection with minimal performance loss. Building on this, [45] employed the last layer of a DNN as a proxy model in their benchmark. Other efforts to improve efficiency revolve around sub-sampling strategies with warm starts [7]. In contrast, rather than improving efficiency of the AL cycle, we update the DNN while constructing the batch, immediately incorporating label information that guides the construction process.
@@ 3 Fast Bayesian Updates for Deep Neural Networks
In this section, we first introduce the general concept of Bayesian updates together with the variant MC-based updates [20, 40]. Afterward, we propose our novel method focusing on an efficient update of the Gaussian posterior distribution via last-layer LAs. For an introduction to LA, we refer to [8].
@@ 3.1 Bayesian Updates
We focus on classification problems with instance space $\mathcal{X}$ and label space $\mathcal{Y}=\{0,\dots,K-1\}$. The primary goal in our setting is to efficiently incorporate the information of new instance-label pairs $\mathcal{D}^{\oplus}=\{({\boldsymbol{x}}_{n},y_{n})\}_{n=1}^{N}\subset\mathcal{X}\times\mathcal{Y}$} into a DNN trained on dataset $\mathcal{D}\subset\mathcal{X}\times\mathcal{Y}$. Retraining the entire network on the extended dataset $\mathcal{D}\cup\mathcal{D}^{\oplus}$ results in high computational cost for a large dataset $\mathcal{D}$. Conversely, using the new data solely can cause catastrophic forgetting [34].
For this purpose, we employ BNNs with Bayesian updates as an efficient alternative to retraining. The main idea of BNNs is to estimate the posterior distribution $p(\boldsymbol{\omega}|\mathcal{D})$ over the parameters $\boldsymbol{\omega}\in\Omega$ given the observed training data $\mathcal{D}$ using Bayes’ theorem. The obtained posterior distribution over the parameters can then be used to specify the predictive distribution over a new instance’s class membership via marginalization:
@@
Thereby, the likelihood $p(y|{\boldsymbol{x}},\boldsymbol{\omega})=[\mathrm{softmax}(f^{\boldsymbol{\omega}}({\boldsymbol{x}}))]_{y}$ denotes the probabilistic output of a DNN with parameters $\boldsymbol{\omega}$, where $f^{\boldsymbol{\omega}}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ is a function outputting class-wise logits. [^1] We denote the $i$-th element of a vector $\boldsymbol{b}$} as ${\boldsymbol{b}}]_{i}=b_{i}$.
The formulation in Eq.˜1 provides a theoretically sound way to obtain updated predictions. In particular, this is because the probabilistic outputs $p(y|{\boldsymbol{x}},\boldsymbol{\omega})$ do not directly depend on the training data $\mathcal{D}$. Consequently, to obtain an updated predictive distribution, we do not need to update the parameters $\boldsymbol{\omega}$ directly but only the posterior distribution $p(\boldsymbol{\omega}|\mathcal{D})$. The updated posterior distribution $p(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ is found through Bayes’ theorem, where the current posterior distribution $p(\boldsymbol{\omega}|\mathcal{D})$ is considered the prior and multiplied with the likelihood $p(y|{\boldsymbol{x}},\boldsymbol{\omega})$ per instance-label pair $({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}$. As instances in $\mathcal{D}$ and $\mathcal{D}^{\oplus}$ are assumed to be independently distributed, we can simplify the likelihood and reformulate the parameter distribution as follows [^2] We denote $p(y_{1},\dots,y_{N}\mid{\boldsymbol{x}}_{1},\dots,{\boldsymbol{x}}_{N},\boldsymbol{\omega})$ with $\mathcal{D}=\{({\boldsymbol{x}}_{n},y_{n})\}_{n=1}^{N}$ as $p(\mathcal{D}|\boldsymbol{\omega})$.:
$p(\boldsymbol{\omega}|\mathcal{D}^{\oplus},\mathcal{D})\propto p(\boldsymbol{\omega}|\mathcal{D})p(\mathcal{D}^{\oplus}|\mathcal{D},\boldsymbol{\omega})\overset{\mathrm{i.i.d.}}{=}p(\boldsymbol{\omega}|\mathcal{D})p(\mathcal{D}^{\oplus}|\boldsymbol{\omega})=p(\boldsymbol{\omega}|\mathcal{D})\prod_{\mathclap{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}}p(y|{\boldsymbol{x}},\boldsymbol{\omega}).$ (2)
We refer to Eq.˜2 as the Bayesian update.
@@
The most common realization [20, 40] of this update is through MC-based BNNs, such as MC-Dropout and deep ensembles. These BNNs rely on samples (or hypotheses) $\boldsymbol{\omega}_{1},\dots,\boldsymbol{\omega}_{M}$ drawn from an approximate posterior $q(\boldsymbol{\omega}|\mathcal{D})$. Research [40, 20] assumes that all hypotheses are equally likely to explain the observed data and have the same probability before updating. By updating the posterior distribution through Eq.˜2, they weigh more likely hypotheses given the new data higher. We refer to these as MC-based updates with a formal definition given in Appendix˜A. Figure˜1 illustrates this concept where different hypotheses $\boldsymbol{\omega}_{1},\dots,\boldsymbol{\omega}_{M}\sim q(\boldsymbol{\omega}|\mathcal{D})$ are shown. Each hypothesis represents a possible true solution for the learning task (white instances). When new data (green instances) arrives, we weigh each hypothesis by its likelihood of explaining the new data and obtain an updated prediction without retraining. This results in an updated predictive distribution, as seen in bold in Fig.˜1 (right).